Find the electric flux in a sphere of radius 0.48 cm from a charge of 55 nC.




1
Expert's answer
2022-05-12T14:04:01-0400
"\phi=\oint EdA =\frac{q}{\epsilon_0}"

"\phi=\frac{55*10^{-9}}{8.85*10^{-12}}=6215\:\rm N \cdot m^2/C"